OpenStudy (anonymous):

Can anyone factor this? (x^2 + 2x + 1)/(x + 1)

OpenStudy (bahrom7893):

(x+1)^2/(x+1)=x+1

OpenStudy (bahrom7893):

the top is of the form: a^2+2ab+b^2, where a=x and b=1

OpenStudy (anonymous):

ah, right! thanks

OpenStudy (andriod09):

\[(x^{2}+2x+1)\div(x+1)\] That correct?